Fatimah:

Which simplified ratio correctly compares 12 yards to 20 feet? 5:3 3:5 5:9 9:5

Fatimah:

The ratio of cheese sandwiches to egg sandwiches on a store shelf is 8 to 18. Simplify the ratio. Enter your answer by filling in the boxes. to

Fatimah:

q2: There are 9 children and 13 adults on a bus. What is the ratio of the total number of people on the bus to the number of children on the bus? Enter your answer by filling in the boxes. to

Fatimah:

There are 9 children and 13 adults on a bus. What is the ratio of the total number of people on the bus to the number of children on the bus? Enter your answer by filling in the boxes. to

adam321:

im sure is 9:5 bcz 12 yard = 36 feet so 36:12 (4*9=36) (4*5=20)